Consider the functions f(x) = sqrt x-1 and g(x) = ax+b. In the standard (x,y) coordinate plane, y = f(g(x)) passes through (0,1) and (2,3). What is the value of a+b?
So far I have f(g(x)) = sqrt (ax+b) -1, slope = 2, y-3 = sqrt(2x-5). Now I don't know what to do or if I am even doing it correctly.

Answer by rothauserc(4718)

f(g(x)) = square root(ax + b) - 1
:
1) square root(a*0+b) -1 = 1
square root(b) = 2
b = 4
:
2) square root(a*2+4) -1 = 3
square root(2a+4) = 4
2a+4 = 16
a = 6
:
**************
a+b = 6+4 = 10
**************
: